A project yields an annual benefit of $25 a year, starting next year and continuing forever. What is the present value of the benefits if the interest rate is 10%?
Answer #1

Answer

the project is a forever series.

The present value of forever series =A/i

A=Annual benefit

i=interest rate

PV=25/0.1=$250

The present value of the series is $250